OpenStudy (anonymous):

Solve the equation 2/5x − 3 = 13 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

Just use the cross-multiplication method.

OpenStudy (jiteshmeghwal9):

\[\LARGE{\frac{2}{5x-3}=\frac{13}{1} }\]by cross multiplication\[\LARGE{2=13(5x-3)}\]solve this from here :)

OpenStudy (anonymous):

thx

OpenStudy (jiteshmeghwal9):

yw :)